OpenStudy (anonymous):

What is the result of adding the system of equations? 2x + y = 4 3x - y = 6

OpenStudy (ivettef365):

just add 2x +y = 4 3x -y = 6 ------------

OpenStudy (anonymous):

5x=10?

OpenStudy (anonymous):

right @santisanchez

OpenStudy (ivettef365):

correct, now divide by 5 both sides to find value of x

OpenStudy (uri):

2x +y = 4 3x -y = 6 ------------ 5x =10 x =2

OpenStudy (anonymous):

thank you guys!!

OpenStudy (uri):

yw!

OpenStudy (ivettef365):

when you have value of x then you need to find value of y

OpenStudy (uri):

Take the second eq or first eq your choice and put 2 instead of x Let's take second eq 3x -y = 6 3(2) -y =6 6-y=6 -y =6/6 y=-1?